# YITP-25-79, IPMU25-0028, RIKEN-iTHEMS-Report-25

### v1

@@ Abstract
b Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study (UTIAS), The University of Tokyo, Kashiwa, Chiba 277-8583, Japan
c The Hakubi Center for Advanced Research, Kyoto University, Yoshida Ushinomiyacho, Sakyo-ku, Kyoto 606-8501, Japan
d RIKEN iTHEMS, Wako, Saitama, 351-0198, Japan
e Department of Physics, College of Humanities and Sciences, Nihon University, Tokyo 156-8550, Japan
f High Energy Physics Research Unit, Department of Physics, Faculty of Science, Chulalongkorn University, Pathumwan, Bangkok 10330, Thailand
1 Introduction
@@
In the late stages of inspiral, compact objects such as black holes (BHs) can be deformed by external tidal fields. These deformations encode information about the properties of compact objects and the underlying theory of gravity, and are captured by their tidal response functions. The leading conservative tidal effects are quantified by the so-called tidal Love numbers (TLNs) [1, 2, 3], which characterize how easily an object deforms in response to an external tidal field. TLNs give rise to phase corrections to the emitted GWs, starting at 5 post-Newtonian (5PN) order  [4, 5, 6]. These effects were constrained by using data from the GW event GW170817 [7, 8, 9]. In four-dimensional general relativity (GR), it was shown that the vacuum solutions, e.g., Kerr BHs, have vanishing TLNs for generic multipole modes with $\ell\geq 2$ [2, 3, 10, 11, 12, 13, 14, 15]. However, BHs in theories beyond four-dimensional GR generally can have non-vanishing TLNs [16, 17, 18, 19, 20, 21, 22, 23]. In addition, horizonless compact objects that may arise due to quantum gravity corrections [24, 16, 25, 26, 27, 28, 29] or BHs surrounded by matter fields in general possess non-zero TLNs [30, 31, 32, 33, 34]. In fact, TLNs can also exhibit a logarithmic scale dependence, indicating “running” TLNs (see, e.g., [30, 12, 21, 32, 34, 23] and references therein). The detection of non-vanishing TLNs of BHs would provide compelling evidence for deviations from GR in vacuum. These deviations can be analyzed within a parametrized formalism [34], as long as they are described by a decoupled master equation with a single degree of freedom. [^1] See [35, 36] for the parametrized BH quasinormal ringdown formalism.
@@
In addition to the conservative tidal effects, there are dissipative tidal responses that lead to energy loss and affect the phase of gravitational waveforms. These effects enter at 4PN order for non-rotating objects and at 2.5PN order for spinning objects [37, 38, 39, 40]. These dissipative effects are quantified by the so-called tidal dissipation numbers (TDNs) [41, 33, 42]. In the case of BHs, the dissipation process is attributed to the absorption of gravitational perturbations by the event horizon. [^2] In the case of neutron stars, TDNs account for energy loss due to viscous dissipation (bulk/shear viscosity) of the internal fluids (see [43, 44, 45] and references therein). Thus, measuring TDNs allows us to observationally distinguish BHs from horizonless compact objects that may arise from Planckian corrections [46, 25, 27, 47, 48, 29]. TDNs were calculated both within four-dimensional GR [13, 15, 41, 49, 50, 51, 29, 52, 33, 53, 54] and extensions beyond GR [42]. It is important to note that in extreme mass ratio inspirals, even small deviations in TDNs can accumulate over thousands of orbital cycles, leaving detectable imprints on the gravitational waveform.
@@
Unlike the case of TLNs, a parametrized BH TDN formalism has not yet been developed in the literature. In this paper, we establish for the first time such a parametrized framework for BH TDNs, revealing their general properties in a model-agnostic way. Our approach is based on introducing perturbative modifications to the master equations—namely, the Regge-Wheeler/Zerilli equations [55, 56]—which describe the dynamics of metric perturbations around a static and spherically symmetric BH. In order to analytically solve the perturbation equations, we use the Mano-Suzuki-Takasugi (MST) method [57, 58, 59], and examine how small deviations in the effective potential change the resulting physical observables. By performing the asymptotic matching between the MST horizon-ingoing solution and the far-zone perturbations obtained via the worldline EFT approach (see, e.g., [60, 12, 15]), we unambiguously extract the static TLNs and TDNs. In this framework, we use the so-called renormalized angular momentum to clearly distinguish between the tidal source and the induced multipole moment. This procedure ensures that the static TLNs and TDNs are uniquely determined, and remain unaffected by ${\cal O}(1/r)$ corrections arising from the redefinition of the master variable [35, 22].
The rest of the paper is organized as follows. In Section 2, we briefly review the relativistic formulation of TDNs in GR based on the MST method. In Section 3, we establish the parametrized BH TDN formalism. In Section 4, we apply our parametrized formalism to compute TDNs in particular theories. we draw our conclusions in Section 5. Throughout this paper, we use the unit $c=G=1$.
@@ 2.2 Tidal response from MST method
with $\lambda\equiv\ell(\ell+1)-2$. We use the superscript “$+$” to denote quantities defined in the even-parity sector, whereas “--” stands for those in the odd-parity sector. Note that the radial equations do not depend on the azimuthal magnetic quantum number $m$ due to the spherical symmetry of the background.
@@ 2.2.1 Near-zone solution in the odd-parity sector
Note also that, if $\nu$ is a solution of Eq. (2.9), then $-\nu-1$ is also a solution. We see from Eq. (2.13) that the decaying mode $\psi_{-\nu-1}^{(0)}(x)$, defined in Eq. (2.11), becomes
@@ 2.2.2 Far-zone (radiative) solution
×$\quad\times\sum_{n=-\infty}^{\infty}a_{n}^{\nu}\frac{\Gamma(g)\Gamma(g+2)}{\Gamma(g-h+1)\Gamma(-h+1)}(2i\varepsilon x)^{n}\,{}_{1}F_{1}(-h-1,g-h+1;2i\varepsilon x)\;.$ (2.16)
@@ 2.2.3 Tidal response for odd-parity sector
${\cal F}_{\nu}^{-,\,{\rm Bare}}(\omega)\equiv\frac{\mathopen{}\mathclose{{}\left(\text{the coefficient of $x^{-\nu}$ in $\psi_{-\nu-1}^{(0)}(x)$}}\right)}{\mathopen{}\mathclose{{}\left(\text{the coefficient of $x^{\nu+1}$ in $\psi_{\nu}^{(0)}(x)$}}\right)}\;.$ (2.22)
@@ 2.3 Chandrasekhar transformation
±${\cal H}_{\ell}^{\pm}\equiv\mathopen{}\mathclose{{}\left(f\frac{{\rm d}}{{\rm d}r}}\right)^{2}+\omega^{2}-fV_{\ell}^{\pm}\;,$ (2.29)
which allow us to write the Zerilli/Regge-Wheeler equations as ±${\cal H}_{\ell}^{\pm}\psi_{\ell}^{\pm}=0$. The Chandrasekhar transformation is defined by the following operators:
@@
Recall that $\lambda=(\ell+2)(\ell-1)$ and $x=r/r_{g}$. Note that the operators ${\cal H}_{\ell}^{\pm}$ can be written into the form,
±${\cal H}_{\ell}^{\pm}=\frac{1}{r_{g}^{2}}\Bigg{[}{\cal D}_{\ell}^{\pm}{\cal D}_{\ell}^{\mp}+\mathopen{}\mathclose{{}\left(\frac{\lambda^{2}(\lambda+2)^{2}}{36}+\varepsilon^{2}}\right)\Bigg{]}\;.$ (2.31)
Then, one immediately finds that ±${\cal H}_{\ell}^{\pm}{\cal D}_{\ell}^{\pm}={\cal D}_{\ell}^{\pm}{\cal H}_{\ell}^{\mp}$, and therefore $\psi_{\ell}^{\pm}\propto{\cal D}_{\ell}^{\pm}\psi_{\ell}^{\mp}$. Note that the Chandrasekhar transformation preserves the asymptotic behaviors of the mode functions up to an overall constant factor. Thus, in vacuum GR, the odd- and even-parity sectors share the same static TLNs and TDNs. For the TDNs, we write
@@ 3 Parametrized TDN formalism
Note that the factor of $r_{g}^{-2}$ has been inserted in front of $\delta V_{\ell}^{\pm}$ in Eq. (3.2), so that $\delta V_{\ell}^{\pm}$ is dimensionless. Since the term with $\omega^{2}$ is irrelevant for the TDNs, we will neglect this term. Then, all the deviations from GR are absorbed in $\delta V_{\ell}^{\pm}$, which we parametrize as a series expansion in powers of $r_{g}/r$ as follows:999 Note that $\delta V_{\ell}^{\pm}$ in Eq. (3.4) does not involve terms with $j<3$. Otherwise, an infrared cutoff must be introduced to regularize some of the integrals in the computations that follow.
@@ 3.1 Odd-parity sector
$\begin{split}\delta{\cal F}_{\nu}^{-,\,{\rm Bare}}&=\Big{(}\delta C_{+}^{-}+\delta C_{-}^{-}\Big{)}\times\frac{\mathopen{}\mathclose{{}\left(\text{the coefficient of $x^{-\nu}$ in $\psi_{-\nu-1}^{(0)}(x)$}}\right)}{\mathopen{}\mathclose{{}\left(\text{the coefficient of $x^{\nu+1}$ in $\psi_{\nu}^{(0)}(x)$}}\right)}\;.\end{split}$ (3.12)
@@
×$\quad\times\sum_{j\geq 3}\alpha_{j}^{-}\bigg{\{}\mathopen{}\mathclose{{}\left[1+2i\varepsilon\mathopen{}\mathclose{{}\left(\ell_{2}-\frac{\ell^{2}+\ell+4}{2\ell(\ell+1)}}\right)}\right]e^{-,\,j}_{(0)}+i\varepsilon{\cal N}_{\ell}^{-}e^{-,\,j}_{(1)}\bigg{\}}\;.$ (3.16)
@@
×$\quad\times\sum_{j\geq 3}\alpha_{j}^{-}\bigg{\{}\mathopen{}\mathclose{{}\left[1+2i\varepsilon\mathopen{}\mathclose{{}\left(\ell_{2}-\frac{\ell^{2}+\ell+4}{2\ell(\ell+1)}}\right)}\right]e^{-,\,j}_{(0)}+i\varepsilon{\cal N}_{\ell}^{-}e^{-,\,j}_{(1)}\bigg{\}}$
@@
[…]
@@
[…]
@@ 3.2 Even-parity sector
[…]
@@
The numerical values of $e^{+,\,j}_{(0)}$ for several values of $j$ can be found in [34]. We note that for positive $\alpha_{j}^{+}$ (i.e., when the effective potential receives positive corrections), the TDNs receive negative corrections since the basis $e^{+,\,j}_{(1)}$ in Table 3 are all negative [see Eq. (3.24)]. Hence, in this particular case, the TDNs are lower than the GR value.
@@ 4.2 Einstein-Maxwell system
for $\ell=2,3,4$, respectively. Note that in our formalism based on the MST method, the coefficient of $x^{-\ell}$ in the first-order solution does not directly correspond to the TDNs. This, in fact, leads to differences between our results and those obtained in [54].

### v2

@@ Abstract
b Research Center for the Early Universe (RESCEU), Graduate School of Science, The University of Tokyo,
Hongo 7-3-1, Bunkyo-ku, Tokyo 113-0033, Japan
c Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study (UTIAS), The University of Tokyo, Kashiwa, Chiba 277-8583, Japan
d The Hakubi Center for Advanced Research, Kyoto University, Yoshida Ushinomiyacho, Sakyo-ku, Kyoto 606-8501, Japan
e RIKEN iTHEMS, Wako, Saitama, 351-0198, Japan
f Department of Physics, College of Humanities and Sciences, Nihon University, Tokyo 156-8550, Japan
g High Energy Physics Research Unit, Department of Physics, Faculty of Science, Chulalongkorn University, Pathumwan, Bangkok 10330, Thailand
1 Introduction and summary
@@
In the late stages of inspiral, compact objects such as black holes (BHs) can be deformed by external tidal fields. These deformations encode information about the properties of compact objects and the underlying theory of gravity, and are captured by their tidal response functions. The leading conservative tidal effects are quantified by the so-called tidal Love numbers (TLNs) [1, 2, 3], which characterize how easily an object deforms in response to an external tidal field. TLNs give rise to phase corrections to the emitted GWs, starting at 5 post-Newtonian (5PN) order  [4, 5, 6]. These effects were constrained by using data from the GW event GW170817 [7, 8, 9]. In four-dimensional general relativity (GR), it was shown that the vacuum solutions, e.g., Kerr BHs, have vanishing TLNs for generic multipole modes with $\ell\geq 2$ [2, 3, 10, 11, 12, 13, 14, 15]. However, BHs in theories beyond four-dimensional GR generally can have non-vanishing TLNs [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. In addition, horizonless compact objects that may arise due to quantum gravity corrections (see, e.g., [26, 16, 27, 28, 29, 30, 31] and references therein) or BHs surrounded by matter fields in general possess non-zero TLNs [32, 33, 34, 35, 22]. In fact, TLNs can also exhibit a logarithmic scale dependence, indicating “running” TLNs (see, e.g., [32, 12, 21, 34, 22, 24] and references therein). The detection of non-vanishing TLNs of BHs would provide compelling evidence for deviations from GR in vacuum. These deviations can be analyzed within a parametrized formalism [22], as long as they are described by a decoupled master equation with a single degree of freedom. [^1] See [36, 37] for the parametrized BH quasinormal ringdown formalism, which applies to Regge-Wheeler/Zerilli equations with small corrections. See also [38] for an extension of the formalism to modified Teukolsky equations [39, 40].
@@
In addition to the conservative tidal effects, there are dissipative tidal responses that lead to energy loss and affect the phase of gravitational waveforms. These effects enter at 4PN order for non-rotating objects and at 2.5PN order for spinning objects [41, 42, 43, 44]. These dissipative effects are quantified by the so-called tidal dissipation numbers (TDNs) [45, 35, 46]. In the case of BHs, the dissipation process is attributed to the absorption of gravitational perturbations by the event horizon. [^2] In the case of neutron stars, TDNs account for energy loss due to viscous dissipation (bulk/shear viscosity) of the internal fluids (see [47, 48, 49] and references therein). Thus, measuring TDNs allows us to observationally distinguish BHs from horizonless compact objects that may arise from Planckian corrections [50, 27, 29, 51, 52, 53, 31]. TDNs were calculated both within four-dimensional GR [13, 15, 45, 54, 55, 56, 31, 57, 35, 58, 59] and extensions beyond GR [46]. It is important to note that in extreme mass ratio inspirals, even small deviations in TDNs can accumulate over thousands of orbital cycles, leaving detectable imprints on the gravitational waveform (see e.g., [51, 60, 52, 53, 61]). In the case of comparable-mass binaries, the importance of tidal dissipation effects in distinguishing between BHs and neutron stars has been pointed out [62, 63], and the measurability in the current and next-generation observations is discussed in [64, 65, 66].
@@
Unlike the case of TLNs, a parametrized BH TDN formalism has not yet been developed in the literature. In this paper, we establish for the first time such a parametrized framework for BH TDNs, revealing their general properties in a model-agnostic way. Our approach is based on introducing perturbative modifications to the master equations—namely, the Regge-Wheeler/Zerilli equations [67, 68]—which describe the dynamics of metric perturbations around a static and spherically symmetric BH. In order to analytically solve the perturbation equations, we use the Mano-Suzuki-Takasugi (MST) method [69, 70, 71], and examine how small deviations in the effective potential change the resulting physical observables. By performing the asymptotic matching between the MST horizon-ingoing solution and the far-zone perturbations obtained via the worldline effective field theory (EFT) approach (see, e.g., [72, 12, 15]), we unambiguously extract the static TLNs and TDNs. In this framework, we use the so-called renormalized angular momentum to clearly distinguish between the tidal source and the induced multipole moment. This procedure ensures that the static TLNs and TDNs are uniquely determined, and remain unaffected by ${\cal O}(1/r)$ corrections arising from the redefinition of the master variable [36, 23].
Outline and summary
Here, we summarize the outline and results of our paper. Our main results are given by Eqs. (3.18) and (3.23) together with those shown in Tables 1, 2, and 3. These are obtained via the following step-by-step workflow:
  1. (i)
    In Section 2, we review the definition of the tidal response function [Eq. (2.22)] from the asymptotic behaviors of the near-zone solution using the MST method. Based on this definition, we can unambiguously extract the TDNs of a Schwarzschild BH in the odd-parity sector as in Eq. (2.25).
  2. (ii)
    In Section 3, we first define our parametrized TDN formalism based on the deformed Zerilli/Regge-Wheeler equations (3.2). Here, deviations from the vacuum GR, appearing in the modifications of the effective potential (3.4), are controlled by small coefficients $\alpha_{j}^{\pm}$ in the $r_{g}/r$-expansion, where $r_{g}$ is the horizon radius and the superscript $\pm$ corresponds to the parity even/odd.
  3. (iii)
    We then solve Eq. (3.2) up to first order in $\alpha_{j}^{\pm}$ and in the frequency $\omega$, using the Green’s function method. By using the vacuum-GR Green’s function [Eq. (3.9)], the first-order solutions in the odd-parity sector are formally given by Eq. (3.10).
  4. (iv)
    Using the asymptotic behaviors of the first-order solutions [Eq. (3.11)], we define the modification of the response function as in Eq. (3.12). This can be written in terms of the bases $e^{-,\,j}_{(0)}$ and $e^{-,\,j}_{(1)}$ defined in Eq. (3.15). Analytic expressions and numerical values of $e^{-,\,j}_{(1)}$ for $\ell=2,3,4$ are presented in Tables 1 (with $j\geq 3$) and 2 (with $3\leq j\leq 9$), respectively.
  5. (v)
    Finally, from Eq. (3.17), we read off the static TLNs and the TDNs for the odd-parity sector as in Eq. (3.18). A similar methodology can be straightforwardly applied to the even-parity sector, resulting in Eq. (3.23). Numerical values of $e^{+,\,j}_{(1)}$ for $\ell=2,3,4$ with $3\leq j\leq 9$ are shown in Table 3.
After obtaining our main results, in Section 4, we apply our formalism to compute odd-parity TDNs in three examples: the EFT of BH perturbations with timelike scalar profile [73, 74], the Einstein-Maxwell system, and the higher curvature-extension of GR [75, 17]. Finally, we draw our conclusions in Section 5. Throughout this paper, we use the unit $c=G=1$.
@@ 2.2 Tidal response from MST method
with $\lambda\equiv\ell(\ell+1)-2$. We use the superscript “++” to denote quantities defined in the even-parity sector, whereas “-” stands for those in the odd-parity sector. Note that the radial equations do not depend on the azimuthal magnetic quantum number $m$ due to the spherical symmetry of the background.
@@ 2.2.1 Near-zone solution in the odd-parity sector
Note also that, if $\nu$ is a solution of Eq. (2.9), then $\nu-1$ is also a solution. We see from Eq. (2.13) that the decaying mode $\psi_{-\nu-1}^{(0)}(x)$, defined in Eq. (2.11), becomes
@@ 2.2.2 Far-zone (radiative) solution
$\quad\times\sum_{n=-\infty}^{\infty}a_{n}^{\nu}\frac{\Gamma(g)\Gamma(g+2)}{\Gamma(g-h+1)\Gamma(-h+1)}(2i\varepsilon x)^{n}\,{}_{1}F_{1}(-h-1,g-h+1;2i\varepsilon x)\;.$ (2.16)
@@ 2.2.3 Tidal response for odd-parity sector
[…]
@@ 2.3 Chandrasekhar transformation
${\cal H}_{\ell}^{\pm}\equiv\mathopen{}\mathclose{{\left(f\frac{{\rm d}}{{\rm d}r}}}\right)^{2}+\omega^{2}-fV_{\ell}^{\pm}\;,$ (2.29)
which allow us to write the Zerilli/Regge-Wheeler equations as ${\cal H}_{\ell}^{\pm}\psi_{\ell}^{\pm}=0$. The Chandrasekhar transformation is defined by the following operators:
@@
[…]
${\cal H}_{\ell}^{\pm}=\frac{1}{r_{g}^{2}}\Bigg[{\cal D}_{\ell}^{\pm}{\cal D}_{\ell}^{\mp}+\mathopen{}\mathclose{{\left(\frac{\lambda^{2}(\lambda+2)^{2}}{36}+\varepsilon^{2}}}\right)\Bigg]\;.$ (2.31)
Then, one immediately finds that ${\cal H}_{\ell}^{\pm}{\cal D}_{\ell}^{\pm}={\cal D}_{\ell}^{\pm}{\cal H}_{\ell}^{\mp}$, and therefore $\psi_{\ell}^{\pm}\propto{\cal D}_{\ell}^{\pm}\psi_{\ell}^{\mp}$. Note that the Chandrasekhar transformation preserves the asymptotic behaviors of the mode functions up to an overall constant factor. Thus, in vacuum GR, the odd- and even-parity sectors share the same static TLNs and TDNs. For the TDNs, we write
@@ 3 Parametrized TDN formalism
Note that the factor of $r_{g}^{-2}$ has been inserted in front of $\delta V_{\ell}^{\pm}$ in Eq. (3.2), so that $\delta V_{\ell}^{\pm}$ is dimensionless. Since the term with $\omega^{2}$ is irrelevant for the TDNs, we will neglect this term. Then, all the deviations from GR are absorbed in $\delta V_{\ell}^{\pm}$, which we parametrize as a series expansion in powers of $r_{g}/r$ as follows: [^9] Note that $\delta V_{\ell}^{\pm}$ in Eq. (3.4) does not involve terms with $j<3$. Otherwise, an infrared cutoff must be introduced to regularize some of the integrals in the computations that follow.
@@ 3.1 Odd-parity sector
[…]
@@
$\quad\times\sum_{j\geq 3}\alpha_{j}^{-}\bigg\{\mathopen{}\mathclose{{\left[1+2i\varepsilon\mathopen{}\mathclose{{\left(\ell_{2}-\frac{\ell^{2}+\ell+4}{2\ell(\ell+1)}}}\right)}}\right]e^{-,\,j}_{(0)}+i\varepsilon{\cal N}_{\ell}^{-}e^{-,\,j}_{(1)}\bigg\}\;.$ (3.16)
@@
$\quad\times\sum_{j\geq 3}\alpha_{j}^{-}\bigg\{\mathopen{}\mathclose{{\left[1+2i\varepsilon\mathopen{}\mathclose{{\left(\ell_{2}-\frac{\ell^{2}+\ell+4}{2\ell(\ell+1)}}}\right)}}\right]e^{-,\,j}_{(0)}+i\varepsilon{\cal N}_{\ell}^{-}e^{-,\,j}_{(1)}\bigg\}$
@@
$\lim_{j\to\mathbb{Z}_{\geq 3}}\bigg\{\thinspace$ $\frac{1}{2(j-3)}+\frac{16}{15(j-4)}+\frac{3}{j-5}+\frac{14}{j-6}-\frac{120}{j-8}+\frac{108}{j-9}$
@@
$\lim_{j\to\mathbb{Z}_{\geq 3}}\bigg\{\thinspace$ $\frac{1}{2(j-3)}+\frac{8}{5(j-4)}+\frac{113}{15(j-5)}+\frac{318}{5(j-6)}-\frac{2520}{j-8}+\frac{7812}{j-9}-\frac{8624}{j-10}+\frac{9800}{3(j-11)}$
@@ 3.2 Even-parity sector
Since the analytic expressions of these bases are complicated, we directly obtain their numerical values. In Table 3, we summarize the numerical values of $e^{+,\,j}_{(1)}$ for $\ell=2,3,4$ with $3\leq j\leq 9$. For completeness, we express the static TLNs and the TDNs for the even-parity sector as
@@
For $\ell=2,3,4$, the numerical values of $e^{+,\,j}_{(0)}$ for $j\geq 2\ell+4$ coincide with the results in [22] by multiplying the normalization factor for each $\ell$. [^12] The mismatch at small $j$ is explained in Appendix B of [23] for the odd-parity case. In addition, one should take care of the logarithmic dependence $\log\mathopen{}\mathclose{{\left(r/r_{g}}}\right)$ for $j\leq 2\ell+3$ for the basis of the TLNs. We note that for positive $\alpha_{j}^{+}$ (i.e., when the effective potential receives positive corrections), the TDNs receive negative corrections since the basis $e^{+,\,j}_{(1)}$ in Table 3 are all negative [see Eq. (3.24)]. Hence, in this particular case, the TDNs are smaller than the GR value.
@@ 4.2 Einstein-Maxwell system
for $\ell=2,3,4$, respectively.
Let us comment on the discrepancy between our results and those obtained in [46], which is the same as Appendix B of [23] in the case of static TLNs. Note that in our formalism based on the MST method, the coefficient of $x^{-\ell}$ in the first-order solution does not directly correspond to the TDNs. In our approach, we extract the TDNs using the analytic continuation of $\ell$, which is automatically implemented by the renormalized angular momentum $\nu$. This procedure does not change the total solution itself, but unambiguously separates the source and response contributions. Also, this prescription guarantees that the TDNs we defined remain invariant under redefinition of master variables (see discussion below Eq. (3.33) of [23]). Physically, this involves removing the PN correction for the metric sourced by a point particle and isolating the contribution solely from finite-size effects as TDNs (see, e.g., [82, 45]).